# Factorized Motion Fields for Fast Sparse Input Dynamic View Synthesis

### v1

@@ Abstract.
submissionid: 969journal: TOGccs: Computing methodologies Renderingccs: Computing methodologies Volumetric modelsccs: Computing methodologies Computer visionccs: Computing methodologies Computational photographyccs: Computing methodologies Regularization
@@ 1. Introduction
Synthesizing novel views of a scene given images/videos of the scene in other viewpoints has several applications such as meta-verse, sports, and remote exploration. Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) brought in a seminal shift in novel view synthesis by incorporating differential volume rendering with a compact continuous-depth model. However, it has several limitations such as the inability to handle object motion in the scene as well as long training and rendering times. Recently, K-Planes (Fridovich-Keil et al., 2023) significantly reduced the time complexity for dynamic view synthesis by proposing a factorized 4D volume representation. However, K-Planes requires a large number of input viewpoints to render photo-realistic novel views when employing a multi-view camera setup. The goal of our work is to overcome this limitation and design a fast dynamic radiance field that can effectively learn the dynamic scene with few input viewpoints. We mainly focus on a sparse multi-view camera setting, where a video from each camera viewpoint is available.
Optimizing a dynamic radiance field with few input viewpoints is highly under-constrained. A popular approach to regularize an under-constrained system is to impose additional priors during the optimization. For example, sparse static radiance fields are regularized with depth (Deng et al., 2022; Uy et al., 2023), visibility (Somraj and Soundararajan, 2023) and natural image (Wu et al., 2023a) priors. However, the object motion in dynamic scenes motivates the study of motion priors for sparse input dynamic view synthesis. Since K-Planes employs a 4D volumetric representation without a motion model, it does not allow the motion implicitly learned by the model to be regularized using motion priors. Thus, there is a need to design a dynamic radiance field with an explicit motion field that lends itself to be constrained with motion supervision.
We design a dynamic radiance field consisting of two models, a 5D radiance field that learns the 3D scene at a canonical time instant and a 4D motion field that learns the motion from any time instant $t$ to the canonical time instant $t^{\prime}$. Since the motion model is a one-directional mapping from $t$ to $t^{\prime}$, it is not obvious how to impose flow priors across two arbitrary time instants. We achieve this by constraining our motion model to map a pair of matched points, obtained using the motion prior, to the same 3D point in the canonical volume. Fig. 2 illustrates our approach of imposing the motion prior. This allows us to impose motion priors across any two time instants and across any cameras.
@@
We evaluate our model on two popular multi-view dynamic scene datasets and find that our model outperforms the state-of-the-art dynamic view synthesis models with fewer input viewpoints. We refer to our model as ReFlow-DyRF, since it employs reliable flow priors to regularize the dynamic radiance field. We summarize the main contributions of our work in the following:
@@
  • We achieve very good dynamic view synthesis performance on two popular multi-view datasets with as few as three viewpoints.
@@ 2. Related Work
We provide a quick overview of prior work that share a few attributes of our model in Tab. 1. To the best of our knowledge, ours is the first work to address dynamic view synthesis with few input viewpoints while achieving fast training and rendering.
@@
Dynamic View Synthesis: Dynamic radiance fields can be broadly classified into two categories based on how the temporal modeling is handled, which is crucial in the sparse input setting. A simple approach is to model the dynamic radiance field as a 6D function of position, time, and viewing direction. To utilize flow priors, NSFF (Li et al., 2021) predicts flow as an auxiliary task and constrains the volume density of the mapped 3D points provided by the flow. K-Planes (Fridovich-Keil et al., 2023) and HexPlane (Cao and Johnson, 2023) extend TensoRF (Chen et al., 2022) to a 4D model that maps the position and time to a latent feature, which is then decoded by a tiny multi-layer perceptron (MLP) or using spherical harmonics. Instead of conditioning the radiance field directly on time, DyNeRF (Li et al., 2022) conditions the radiance field on a per-time-instant feature vector which is jointly optimized with the radiance field. The lack of a motion model in these approaches makes it incompatible to impose motion priors when learning with sparse input viewpoints.
@@
The second set of models employs a motion or deformation field that maps the 3D points from a given time instant to a canonical time instant (Pumarola et al., 2021; Wang et al., 2023). TiNeuVox (Fang et al., 2022), SWAGS (Shaw et al., 2023) and CoGS (Yu et al., 2023) replace the scene representation MLP in D-NeRF (Pumarola et al., 2021) with a TensoRF or 3DGS model, but use MLPs to model the motion field, leading to expensive training and rendering time. Different from the above works, prior (Guo et al., 2023) and concurrent works (Wu et al., 2023b) employ a motion model that maps the 3D points from the canonical time instant to a given time instant. However, the forward warping of points can lead to holes in novel views, which may need to be infilled using a separate inpainting network (Guo et al., 2023; Mirzaei et al., 2023; Somraj et al., 2022).
@@
Radiance Fields with Sparse Views: Optimizing a radiance field with sparse input views for static scenes has rich literature. With sparse inputs, the volume rendering equations are severely under-constrained leading to distortions such as floaters and clouds. Few-shot NeRF models address this by constraining the model further with hand-crafted or learned priors. While the early works employ simple priors such as sparsity of mass (Kim et al., 2022) and smoothness of depth (Niemeyer et al., 2022), later works found that the use of richer priors can lead to better performance. The most popular prior is the depth prior (Deng et al., 2022; Roessle et al., 2022; Uy et al., 2023; Somraj et al., 2023), while other priors based on visibility (Somraj and Soundararajan, 2023), semantic information (Jain et al., 2021) and diffusion models (Wynn and Turmukhambetov, 2023; Wu et al., 2023a) are also explored. While most of the works employ the NeRF baseline leading to long training time, recent works explore the use of explicit models (Shi et al., 2023; Xiong et al., 2023; Zhu et al., 2023). However, none of these approaches analyze or regularize the distortions caused by motion in a dynamic scene.
@@
Figure 1. Model architecture: We decompose the dynamic radiance field (DyRF) into a 4D scene flow field $\mathcal{F}_{f}$ that maps a 3D point $\mathbf{p}_{i}$ at time $t$ to the corresponding 3D point $\mathbf{p}_{i}^{\prime}$ at canonical time $t^{\prime}$, and a 5D radiance field $\mathcal{F}_{s}$ that models the scene at canonical time $t^{\prime}$. Both the fields are modeled using a factorized volume (Fridovich-Keil et al., 2023) followed by a tiny MLP, which allows fast optimization and rendering. We note that $\mathbf{G}_{f}$ is modeled using six planes, although we show only three owing to the difficulty in visualizing four dimensions. The MLP $\mathbf{M}_{s}$ is conditioned on time and viewing direction to model time-dependent color variations such as shadows and view-dependent color variations such as specularities. The output of $\mathcal{F}_{s}$ is volume rendered to obtain the color of the pixel and the photometric loss is used to train both the fields. The explicitly modeled motion field $\mathcal{F}_{f}$ is additionally regularized using the flow priors as shown in Fig. 2.
@@
Fast Optimization of Radiance Fields: Fast optimization of radiance fields allows wider usage in real-world applications. Voxel grid based approaches trade memory for speed by replacing the MLP in NeRF with a voxel grid (Fridovich-Keil et al., 2022; Sun et al., 2022). However, the memory requirement of voxel grids increases with the third and fourth power of the grid resolution for static and dynamic scenes respectively. Prior work on fast and compact representation for learning radiance fields can be broadly classified into two categories. One category of models employs a factorized volume representation (Chen et al., 2022; Fridovich-Keil et al., 2023; Cao and Johnson, 2023) to exploit the spatial correlation of the entities to be learned, such as volume density and color. The other category of models (Kerbl et al., 2023; Müller et al., 2022; Barron et al., 2023) mainly exploit the sparsity of the scene to reduce the memory footprint. However, the scene flow for a moving object is non-zero at every time instant, but its variation is temporally correlated. Since the motion field is not sparse temporally, we believe that the sparsification based approaches may tend to learn an independent motion field for every time instant. For example, with a 4DGS model, a single 4D Gaussian may not be able to model the motion across a few time instants and can model the motion at a single time instant only, similar to Lee et al. (2023). On the other hand, factorized volumes can effectively exploit the spatio-temporal correlation of the motion field and hence we employ a 4D factorized model to learn the motion field.
@@ 3.1. Dynamic Radiance Field
We modify the K-Planes (Fridovich-Keil et al., 2023) implementation for dynamic radiance fields to incorporate an explicit motion model. K-Planes employs a multi-resolution grid $\mathbf{G}$ followed by a tiny multi-layer perceptron (MLP) $\mathbf{M}$ to represent a radiance field. As shown in Fig. 1, we model our dynamic radiance field (DyRF) by employing two factorized tensorial models, a radiance field $\mathcal{F}_{s}=\mathbf{G}_{s}\circ\mathbf{M}_{s}$ that represents the 3D scene at a canonical time instant $t^{\prime}$, and a 4D scene flow field $\mathcal{F}_{f}=\mathbf{G}_{f}\circ\mathbf{M}_{f}$ that represents the scene flow for a 3D point $\mathbf{p}$ from time $t$ to $t^{\prime}$. Mapping the 3D scene at every time instant to a canonical volume helps our model enforce temporal consistency of objects in the scene (Pumarola et al., 2021).
@@
We train our DyRF model, similar to K-Planes, by rendering randomly sampled pixels $\mathbf{q}_{t}^{v}$ in view $v$ at time $t$ and using the ground truth color as supervision. To render a pixel $\mathbf{q}_{t}^{v}$, we sample $N_{p}$ points $\left\{\mathbf{p}_{i}\right\}_{i=1}^{N_{p}}$ at depths $\left\{z_{i}\right\}_{i=1}^{N_{p}}$ along the corresponding ray. For every 3D point $\mathbf{p}_{i}$, we first obtain the corresponding 3D point $\mathbf{p}_{i}^{\prime}$ at canonical time $t^{\prime}$ by computing the scene flow from time $t$ to $t^{\prime}$ using $\mathcal{F}_{f}$ as
@@
where $\gamma$ denotes the encoding (Müller et al., 2022) of the viewing direction and the time instant. Conditioning $\mathbf{M}_{s}$ additionally on time allows our model to capture illumination changes due to object motion. The colors of $\mathbf{p}_{i}$ are then volume rendered to obtain the color $\mathbf{c}$ of $\mathbf{q}$ as
@@ 3.2. Training with Flow Priors
2 views K-Planes 17.95 0.62 0.43 0.59 16.51 0.51 0.40 0.35
DyRF 17.98 0.61 0.44 0.67 17.04 0.56 0.37 0.35
ReFlow-DyRF 20.35 0.71 0.33 0.41 19.08 0.66 0.31 0.23
3 views K-Planes 23.65 0.83 0.25 0.34 18.75 0.66 0.30 0.22
DyRF 22.61 0.81 0.27 0.39 18.72 0.68 0.30 0.22
ReFlow-DyRF 25.07 0.88 0.21 0.16 20.41 0.74 0.25 0.12
4 views K-Planes 24.80 0.86 0.20 0.27 23.85 0.86 0.16 0.11
DyRF 23.76 0.83 0.24 0.36 23.51 0.85 0.17 0.13
ReFlow-DyRF 25.86 0.89 0.19 0.17 23.69 0.86 0.17 0.10
Table 2. Quantitative results: We compare our model with K-Planes on N3DV (Li et al., 2022) and InterDigital (Sabater et al., 2017) datasets with two, three, and four input views. We also show the performance of our base DyRF model for reference. We report the PSNR, SSIM, and LPIPS scores for the rendered images and the depth MAE for the rendered depth maps. The best scores in each category are shown in bold.
A dense flow prior across cameras and across time instants is an ideal prior for optimal regularization of our motion field. However, different cameras in a multi-view setting could suffer from variations in parameters such as intrinsics, color balance, exposure and focus making it challenging to determine correspondences across cameras based on luminance constancy. As a result, we find that dense flow estimates obtained using deep optical flow networks are unreliable across cameras as seen in Fig. 3(a). On the other hand, we find that matching keypoints using robust SIFT (Lowe, 2004) descriptors are more reliable across cameras as seen in Fig. 3(b). Specifically, for a pair of frames across time instants $t$ and $s$ and cameras $v$ and $u$, we first extract the SIFT keypoints in individual images using Colmap (Schonberger and Frahm, 2016) and then match the keypoints using SIFT descriptors. By finding the difference between the locations of the matched keypoints, we obtain the sparse flow prior $\mathcal{P}_{\text{sf}}$. While we find that Colmap provides a reasonable number of matches across cameras, newer models such as R2D2 (Revaud et al., 2019) could be employed to improve the richness of the sparse flow prior.
@@ 3.3. Overall Model
Figure 4. Qualitative examples on N3DV dataset with two input views: We can observe that K-Planes finds it hard to learn the moving person leading to significant distortions. Our DyRF model without any priors corrects a few errors by virtue of the common canonical volume. Imposing our priors leads to very good reconstruction.

### v3

@@ Abstract.
submissionid: 969journal: TOGjournalyear: 2024copyright: rightsretainedconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24; July 27-August 1, 2024; Denver, CO, USAbooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24 (SIGGRAPH Conference Papers ’24), July 27-August 1, 2024, Denver, CO, USAdoi: 10.1145/3641519.3657498isbn: 979-8-4007-0525-0/24/07ccs: Computing methodologies Renderingccs: Computing methodologies Volumetric modelsccs: Computing methodologies Computer visionccs: Computing methodologies Computational photographyccs: Computing methodologies Regularization
@@ 1. Introduction
Synthesizing novel views of a scene given images/videos of the scene in other viewpoints has several applications such as meta-verse, sports, and remote exploration. Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) brought in a seminal shift in novel view synthesis by incorporating differential volume rendering with a compact continuous-depth model. However, it has several limitations such as the inability to handle object motion in the scene as well as long training and rendering times. Recently, K-Planes (Fridovich-Keil et al., 2023) significantly reduced the time complexity for dynamic view synthesis by proposing a factorized 4D volume representation. However, K-Planes requires a large number of input viewpoints to render photo-realistic novel views when employing a multi-view camera setup. The goal of our work is to overcome this limitation and design a fast dynamic radiance field that can effectively learn the dynamic scene with few input viewpoints. We mainly focus on a sparse multi view camera setting, where a video from each viewpoint is available.
Optimizing a dynamic radiance field with few input viewpoints is highly under-constrained. A popular approach to regularize an under-constrained system is to impose additional priors during the optimization. For example, sparse static radiance fields are regularized with depth (Deng et al., 2022; Uy et al., 2023), visibility (Somraj and Soundararajan, 2023) and natural image (Wu et al., 2024) priors. However, the object motion in dynamic scenes motivates the study of motion priors for sparse input dynamic view synthesis. Since K-Planes employs a 4D volumetric representation without a motion model, it does not allow the motion implicitly learned by the model to be regularized using motion priors. Thus, there is a need to design a dynamic radiance field with an explicit motion field that lends itself to be constrained with motion supervision.
We design a dynamic radiance field consisting of two models, a 5D radiance field that learns the 3D scene at a canonical time instant and a 4D motion or deformation field that learns the motion from any time instant $t$ to the canonical time instant $t^{\prime}$. Since the motion model is a one-directional mapping from $t$ to $t^{\prime}$, it is not obvious how to impose flow priors across two arbitrary time instants. We achieve this by constraining our motion model to map a pair of matched points, obtained using the motion prior, to the same 3D point in the canonical volume. Fig. 2 illustrates our approach of imposing the motion prior. This allows us to impose motion priors across any two time instants and across any cameras.
@@
We evaluate our model on two popular multi-view dynamic scene datasets and find that our model outperforms the state-of-the-art dynamic view synthesis models with fewer input viewpoints. We refer to our model as RF-DeRF, since we employ Reliable Flow priors for Deformable Radiance Fields. We summarize the main contributions of our work in the following:
@@
  • We achieve very good dynamic view synthesis performance on two popular multi-view datasets with very few views.
@@ 2. Related Work
2.1. Classical Work on Deformation Models
The modeling of dynamic radiance fields through a static radiance field and a motion field is similar to the use of a deformation model that deforms a canonical representation of an entity. This is a popular approach to model deformable solids (Sederberg and Parry, 1986), human motion (Loper et al., 2015), facial expressions (Bradley et al., 2010), deformable garments (Miguel et al., 2012; Pons-Moll et al., 2017), fluids (Adrian, 1991; Gregson et al., 2012; Hawkins et al., 2005), gases (Atcheson et al., 2008), smoke (Hawkins et al., 2005), flames (Ihrke and Magnor, 2004) as well as CT (Computed Tomography) and MRI (Magnetic Resonance Imaging) scans in the medical field (Nunes et al., 2012). Learning the motion or deformation field is also found to improve the reconstruction of 3D scenes (Vedula et al., 2000) and free-viewpoint rendering of dynamic scenes (Carranza et al., 2003).
The canonical space in such deformable models is popularly represented as triangular meshes estimated through a multi-view stereo algorithm (Bradley et al., 2008, 2010; Kraevoy and Sheffer, 2005; Allen et al., 2003), or specialized parametric meshes for humans and faces (Anguelov et al., 2005; Vlasic et al., 2008; Loper et al., 2015), or sums of Gaussians (Stoll et al., 2011). In contrast, we employ a radiance field to learn the canonical space owing to its advantages over meshes (Mildenhall et al., 2020). Different deformation models include free-form deformation (FFD) fields or optical flows (Faloutsos et al., 1997; Wang et al., 2009), bijective mappings (Kraevoy and Sheffer, 2004), morphable models (Blanz and Vetter, 1999), splines (Vlasic et al., 2009), volumetric Laplacian deformations (de Aguiar et al., 2008) and linear systems with basis functions (Ihrke and Magnor, 2004). Our motion field is closer to the free-form deformation among the above deformation approaches, but differs in the representation used for the motion field. Finally, the sparse flow priors we employ can be thought of as similar to tracking the markers (Guskov et al., 2003; Hasler et al., 2006) or keypoints (Bradley et al., 2008; Joo et al., 2018) while estimating the classical deformation models.
2.2. Dynamic View Synthesis
Different from the use of deformable models, Zitnick et al. (2004) learn the 3D scene dynamics using a layered depth representation with motion compensation. More recent volumetric representations such as Multiplane Images (MPI) (Zhou et al., 2018) are extended to handle dynamic scenes by employing an MPI per frame (Lin et al., 2021) or temporal basis functions (Xing and Chen, 2021). However, such approaches suffer from depth discretization artifacts (Mildenhall et al., 2020). Instead of volumetric models, Yoon et al. (2020) employ depth image based warping, where the depth is obtained by combining monodepth with multi-view depth. Recent work on dynamic view synthesis with sparse input viewpoints require depth information obtained through the use of RGB-D cameras (Li et al., 2023) or multi-view stereo (Bansal and Zollhöfer, 2023). Further, such approaches struggle to handle soft edges and translucent objects (Penner and Zhang, 2017).
2.3. Dynamic Radiance Fields
@@
In contrast to the classical approaches, dynamic view synthesis can be solved by learning a 6D radiance field that maps the position, time and viewing direction to the radiance. However, the performance of such approaches degrades when the input viewpoints are sparse. We provide a quick overview of prior work that share a few attributes of our model in Tab. 1. To the best of our knowledge, ours is the first work to address dynamic radiance fields with few input viewpoints while achieving fast training and rendering. In the following, we review the literature on dynamic radiance fields and sparse-input static radiance fields.
@@
Dynamic radiance fields can be broadly classified into two categories based on how the temporal modeling is handled, which is crucial in the sparse input setting. A simple approach is to model the dynamic radiance field as a 6D function of position, time, and viewing direction. To utilize flow priors, NSFF (Li et al., 2021) predicts flow as an auxiliary task and constrains the volume density of the mapped 3D points provided by the flow. K-Planes (Fridovich-Keil et al., 2023) and HexPlane (Cao and Johnson, 2023) extend TensoRF (Chen et al., 2022) to a 4D model that maps the position and time to a latent feature, which is then decoded by a tiny multi-layer perceptron (MLP). Instead of conditioning the radiance field directly on time, DyNeRF (Li et al., 2022) conditions the radiance field on a per-time-instant feature vector which is jointly optimized with the radiance field. The lack of a motion model in these approaches makes it incompatible to impose motion priors when learning with sparse input viewpoints.
@@
The second set of models employs a motion or deformation field that maps the 3D points from a given time instant to a canonical time instant (Pumarola et al., 2021; Wang et al., 2023b, a). TiNeuVox (Fang et al., 2022), SWAGS (Shaw et al., 2023) and CoGS (Yu et al., 2023) replace the scene representation MLP in D-NeRF (Pumarola et al., 2021) with a TensoRF or 3DGS model, but use MLPs to model the motion field, leading to expensive training and rendering time. To achieve fast optimization and rendering, DeVRF (Liu et al., 2022) employs a 4D voxel grid to learn the deformation field, but its memory requirements scale with the fourth power of the grid resolution. Instead of employing a free-form deformation model, Wang et al. (2021) models the motion using discrete cosine transform (DCT) basis functions. Different from the above, prior (Guo et al., 2023) and concurrent work (Wu et al., 2023) employ a motion model that maps the 3D points from the canonical time instant to a given time instant. However, the forward warping of points can lead to holes in novel views, which may need to be infilled using a separate inpainting network (Guo et al., 2023; Mirzaei et al., 2023; Somraj et al., 2022).
2.4. Radiance Fields with Sparse Views
@@
Optimizing a radiance field with sparse input views for static scenes has rich literature. With sparse inputs, the volume rendering equations are severely under-constrained leading to distortions such as floaters. Few-shot NeRF models address this by constraining the model further with hand-crafted or learned priors. While the early works employ simple priors such as sparsity of mass (Kim et al., 2022) and smoothness of depth (Niemeyer et al., 2022), later works found that the use of richer priors can lead to better performance. The most popular prior is the depth prior (Deng et al., 2022; Roessle et al., 2022; Uy et al., 2023; Somraj et al., 2023), while other priors based on visibility (Somraj and Soundararajan, 2023), semantic information (Jain et al., 2021) and diffusion models (Wynn and Turmukhambetov, 2023; Wu et al., 2024) are also explored. While most of the works employ the NeRF baseline leading to long training time, recent works explore the use of explicit models (Shi et al., 2024; Xiong et al., 2023; Zhu et al., 2023). However, none of these approaches analyze or regularize the distortions caused by motion in a dynamic scene.
@@
Figure 1. Model architecture: We decompose the dynamic radiance field into a 4D scene flow or deformation field $\mathcal{F}_{f}$ that maps a 3D point $\mathbf{p}_{i}$ at time $t$ to the corresponding 3D point $\mathbf{p}_{i}^{\prime}$ at canonical time $t^{\prime}$, and a 5D radiance field $\mathcal{F}_{s}$ that models the scene at canonical time $t^{\prime}$. Both the fields are modeled using a factorized volume followed by a tiny MLP, which allows fast optimization and rendering. We note that $\mathbf{G}_{f}$ is modeled using six planes, although we show only three owing to the difficulty in visualizing four dimensions. The MLP $\mathbf{M}_{s}$ is conditioned on time and viewing direction to model time-dependent color variations such as shadows and view-dependent color variations such as specularities. The output of $\mathcal{F}_{s}$ is volume rendered to obtain the color of the pixel and the photometric loss is used to train both the fields. The explicitly modeled motion field $\mathcal{F}_{f}$ is additionally regularized using the flow priors as shown in Fig. 2.
2.5. Fast Optimization of Radiance Fields
@@
Fast optimization of radiance fields allows wider usage in real-world applications. Voxel grid based approaches trade memory for speed by replacing the MLP in NeRF with a voxel grid (Fridovich-Keil et al., 2022; Sun et al., 2022). Prior work on fast and compact representation for learning radiance fields can be broadly classified into two categories. One category of models employs a factorized volume representation (Chen et al., 2022; Fridovich-Keil et al., 2023; Cao and Johnson, 2023) to exploit the spatial correlation of the entities to be learned, such as volume density and color. The other category of models (Kerbl et al., 2023; Müller et al., 2022; Barron et al., 2023) mainly exploit the sparsity of the scene to reduce the memory footprint. However, the scene flow for a moving object is non-zero at every time instant, but its variation is temporally correlated. Since the motion field is not sparse temporally, we believe that the sparsification based approaches may tend to learn an independent motion field for every time instant. For example, with a 4DGS model, a single 4D Gaussian may not be able to model the motion across a few time instants and can model the motion at a single time instant only, similar to Lee et al. (2023). On the other hand, factorized volumes can effectively exploit the spatio-temporal correlation of the motion field and hence we employ a 4D factorized model to learn the motion field.
@@ 3.1. Dynamic Radiance Field
We modify the K-Planes (Fridovich-Keil et al., 2023) implementation for dynamic radiance fields to incorporate an explicit motion model. K-Planes employs a multi-resolution grid $\mathbf{G}$ followed by a tiny multi-layer perceptron (MLP) $\mathbf{M}$ to represent a radiance field. As shown in Fig. 1, we model our dynamic radiance field by employing two factorized tensorial models, a radiance field $\mathcal{F}_{s}=\mathbf{G}_{s}\circ\mathbf{M}_{s}$ that represents the 3D scene at a canonical time instant $t^{\prime}$, and a 4D scene flow or deformation field $\mathcal{F}_{f}=\mathbf{G}_{f}\circ\mathbf{M}_{f}$ that represents the scene flow for a 3D point $\mathbf{p}$ from time $t$ to $t^{\prime}$. Mapping the 3D scene at every time instant to a canonical volume helps our model enforce temporal consistency of objects in the scene (Pumarola et al., 2021).
@@
We train our model, similar to K-Planes, by rendering randomly sampled pixels $\mathbf{q}_{t}^{v}$ in view $v$ at time $t$ and using the ground truth color as supervision. To render a pixel $\mathbf{q}_{t}^{v}$, we sample $N_{p}$ points $\left\{\mathbf{p}_{i}\right\}_{i=1}^{N_{p}}$ at depths $\left\{z_{i}\right\}_{i=1}^{N_{p}}$ along the corresponding ray. For every 3D point $\mathbf{p}_{i}$, we first obtain the corresponding 3D point $\mathbf{p}_{i}^{\prime}$ at canonical time $t^{\prime}$ by computing the scene flow from time $t$ to $t^{\prime}$ using $\mathcal{F}_{f}$ as
@@
where $\gamma$ denotes the encoding (Müller et al., 2022) of the viewing direction and the time instant. Conditioning $\mathbf{M}_{s}$ additionally on time allows our model to capture illumination changes due to object motion. $\mathbf{c}_{i}$ are then volume rendered to obtain the color $\mathbf{c}$ of $\mathbf{q}$ as
@@ 3.2. Training with Flow Priors
2 views HexPlane 13.81 0.41 0.58 2.03 13.78 0.24 0.55 1.51
K-Planes 17.95 0.62 0.43 0.59 16.51 0.51 0.40 0.35
DeRF 17.98 0.61 0.44 0.67 17.04 0.56 0.37 0.35
RF-DeRF 20.35 0.71 0.33 0.41 19.08 0.66 0.31 0.23
3 views HexPlane 15.53 0.49 0.50 1.97 13.85 0.27 0.54 1.43
K-Planes 23.65 0.83 0.25 0.34 18.75 0.66 0.30 0.22
DeRF 22.61 0.81 0.27 0.39 18.72 0.68 0.30 0.22
RF-DeRF 25.07 0.88 0.21 0.16 20.41 0.74 0.25 0.12
4 views HexPlane 15.50 0.47 0.54 1.95 15.18 0.38 0.49 1.24
K-Planes 24.80 0.86 0.20 0.27 23.85 0.86 0.16 0.11
DeRF 23.76 0.83 0.24 0.36 23.51 0.85 0.17 0.13
RF-DeRF 25.86 0.89 0.19 0.17 23.69 0.86 0.17 0.10
Table 2. Quantitative results: We compare our model with K-Planes on N3DV (Li et al., 2022) and InterDigital (Sabater et al., 2017) datasets with two, three, and four input views. We also show the performance of our base DeRF model for reference. We report the PSNR, SSIM, and LPIPS scores for the rendered images and the depth MAE for the rendered depth maps. The best scores in each category are shown in bold.
A dense flow prior across cameras and across time instants is an ideal prior for optimal regularization of our motion field. However, different cameras in a multi-view setting could suffer from variations in parameters such as intrinsics, color balance, exposure and focus making it challenging to determine correspondences across cameras based on luminance constancy. As a result, the dense flow estimates obtained using deep optical flow networks are unreliable across cameras as seen in Fig. 3(a). On the other hand, matching keypoints using robust SIFT (Lowe, 2004) descriptors are more reliable across cameras as seen in Fig. 3(b). Specifically, for a pair of frames across time instants $t$ and $s$ and cameras $v$ and $u$, we extract the SIFT keypoints in individual images using Colmap (Schonberger and Frahm, 2016) and match the keypoints using SIFT descriptors. The difference between the locations of the matched keypoints gives the sparse flow prior $\mathcal{P}_{\text{sf}}$. While Colmap provides a reasonable number of matches, newer models such as R2D2 (Revaud et al., 2019) could be employed to improve the richness of the sparse flow prior.
@@
Figure 4. Qualitative examples on N3DV dataset with two input views: We can observe that K-Planes finds it hard to learn the moving person leading to significant distortions. Our DeRF model (without any priors) corrects a few errors by virtue of the common canonical volume. Imposing our priors leads to much better reconstruction.